# Generalisation Bounds of Zero-Shot Economic Forecasting using Time Series Foundation Models

### v1

@@ 3.2 Model Evaluation
In the forecasting literature, scale-dependent error measures remain fundamental for assessing forecast accuracy. Mean Absolute Error (MAE) and Mean Squared Error/Root Mean Squared Error (MSE/RMSE) in Equations 1 and 2 are the two widely reported scale-dependent accuracy metrics. MAE specifically offers a direct interpretation as the average error in the original units of the series [60].
Both compare a forecast $\hat{y}_{t}$ against the observed value $y_{t}$ over $T$ time steps, but they weigh errors differently. MAE is the arithmetic mean of absolute deviations, squares the residuals before averaging, and then takes the square root. Both MAE and MSE/RMSE compare a forecast $\hat{y}_{t}$ against the observed value $y_{t}$ over $T$ time steps, but their error weighting differs. MAE averages the absolute errors $|y_{t}-\hat{y}_{t}|$, whereas RMSE averages the squared errors $(y_{t}-\hat{y}_{t})^{2}$ before taking the square root.
@@
MAE is easy to interpret the forecast on average.
MSE $=\frac{1}{T}\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2},$ (2)
RMSE $=\sqrt{\text{MSE}}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}}.$
MSE/RMSE as defined in Equation 2 squares errors, giving disproportionate weight to large deviations and making the metric sensitive to outliers or significant irregularities. Both MSE/RMSE and other scale-dependent metrics share a limitation: they are expressed in the original data units, preventing direct comparison across series with different scales. To address this, Symmetric Mean Absolute Percentage Error (SMAPE) in Equation 3 normalises the absolute error by the average magnitude of the actual and forecast values [61],
@@
where $m$ is the seasonal period. This scaling provides a key advantage: MASE is a scale-free error metric. Because the denominator is the MAE of the naive seasonal forecast, the value serves as a direct benchmark. Thus, any MASE $<$ 1 indicates performance superior to the naive baseline, and MASE $>$ 1 denotes inferior accuracy. This scale-free property facilitates widespread adoption in comparative studies of forecasting algorithms. The Diebold-Mariano (DM) in Equation 5, introduced by [63] established a general, loss-function-agnostic framework for testing whether two competing forecasts have the same expected predictive accuracy.
$e_{i,t}=y_{t}-\hat{y}_{i,t}$
be the forecast error from a model $i$, and choose a loss function $\ell(\cdot)$, for example $\ell(e)=e^{2}$ (squared error) or $\ell(e)=|e|$ (absolute error). Define the loss differential
$d_{t}=\ell(e_{1,t})-\ell(e_{2,t}),$
@@
where $\gamma_{k}$ represents the sample autocovariance of $d_{t}$ at lag $k$. Under suitable regularity conditions, $\mathrm{DM}$ is asymptotically distributed as $\mathcal{N}(0,1)$. Consequently, a two-sided $z$-test provides an asymptotic $p$-value. The use of the Newey–West (HAC) estimator in the denominator ensures the test’s validity even when $d_{t}$ follows a moving-average MA($h{-}1$) process, making it applicable interchangeably to MSE/RMSE or MAE comparisons.
@@ 3.3 Zero-Shot Forecasts
TSFMs such as TimeGPT, Chronos, and Moirai that employ zero-shot learning represent a significant advance in predictive analytics. Through large-scale pre-training on a wide variety of time-series datasets, these models develop a generalised understanding of temporal dynamics, capturing regular cycles and anomalous events. Armed with this broad temporal intuition, they can be deployed directly on novel forecasting tasks without requiring extensive domain-specific fine-tuning. At the core of zero-shot forecasting is the idea that, once a foundation model has internalised patterns spanning many industries and time scales, it can transfer that knowledge seamlessly to new contexts. This approach significantly reduces the time, computational resources, and specialised expertise normally needed for forecasting solutions, as it eliminates the need for extensive retraining and adaptation for each new task.
The practical benefits of zero-shot forecasting are most pronounced in settings where conditions change rapidly or data is scarce. For instance, in sectors such as National GDP, Primary Industries, Goods-Producing Industries, and Services Industries, unexpected events and economic shocks can frequently upend historical relationships and data patterns. In such dynamic environments, a zero-shot model can provide immediate, reasonably accurate projections without waiting for new data to accumulate or models to be retrained. Furthermore, the same pre-trained model can often be utilised for both short-term operational decisions and longer-term strategic planning, all without needing to rebuild the model for each specific forecasting horizon.
3.4 Experiment Workflow
@@
We design an experimental back-testing workflow, shown in Figure 2. That compares shift operation persistence and statistical baseline ARIMA models against zero-shot forecasts from TSFMs: TimeGPT-1, Chronos-T5 (small, base, large), and Moirai (small, base, large). We generate one-quarter-ahead forecasts for each model over an expanding training window from 1999Q3 to 2024Q3. The forecasts are produced without any additional fine-tuning; the models were pre-trained before being evaluated on the same error metrics (MAE, RMSE, SMAPE, and MASE). To capture the heterogeneous dynamics of New Zealand’s economy, we apply this experiment to four broad production sectors: National GDP, Primary Industries, Goods-Producing Industries, and Services Industries. The selection of these models for the experiments also reveals which classical and TSFMs are strongest with scarce data.
3.4.1 Persistence model
The persistence forecasting model is given by Equation 6
$\hat{y}*{t+h}=y_{t}\quad\text{for all }h\geq 1,$ (6)
3.4.2 ARIMA model
ARIMA, as implemented by Nixtla [65], is an automatic process that employs a stepwise search procedure to identify optimal ARIMA and seasonal orders. This is achieved by combining unit-root tests for stationarity with the minimization of information criteria. The model acts as a dynamic statistical benchmark, adapting to each sector’s unique autocorrelation structure and seasonality patterns. It applies differencing to handle trend and drift components and fits candidate models using a maximum search space to maintain computational efficiency.
3.4.3 TimeGPT-1 model
TimeGPT-1 is trained on over a billion time series observations drawn from domains such as finance, energy, retail, IoT, and weather, making it the largest publicly documented foundation model for forecasting to date. This extensive training allows the model to generate accurate forecasts on entirely unseen series, such as New Zealand’s GDP and industry sub-series, without any additional fine-tuning or parameter adjustment. Performance is demonstrated by direct comparison to methods like ARIMA and persistence.
3.4.4 Chronos model
Chronos treats time series as a discrete language by scaling and quantizing real-valued observations into tokens. It then trains a transformer model (encoder-decoder or decoder-only) on these tokens using cross-entropy loss. This language-model approach allows Chronos to learn general temporal representations that transfer across datasets, enhancing forecast accuracy in many sectors, and providing a principled way to evaluate how the abstraction of series values compares to models like ARIMA.
3.4.5 Moirai model
Moirai introduces a masked-encoder universal transformer by Salesforce. This pre-trained model is designed for multi-variate, multi-frequency forecasting, capable of capturing long cycles and abrupt shocks within a single framework. And Patch-based input layers learn representations at multiple granularities, enabling the model to handle data from various sectors alongside exogenous covariates without requiring per-task retraining. Moirai uniquely provides flexible predictions, and zero-shot performance, arising from data diversity, allows for direct evaluation against ARIMA and persistence forecasts.
@@ 4.2 Forecast Benchmarking Against State-of-the-art
Figure 5: Comparisons of the Moirai Base Model with LSboost and Factor Models using Reserve Bank Nowcasts.
We extracted forecasts from the RBNZ’s LSboost (Gradient boosting) and Factor models, observing that they are quarterly percentage changes. Consequently, we converted the national GDP data from quarterly annual percentage change to quarterly percentage change to match the forecast format. For comparison in Figure 5, we also used Moirai Base to forecast quarterly percentage change values. The table includes Diebold Mariano two-sided tests computed against an ARIMA model for the period 2009Q1 - 2019Q1, evaluating the RBNZ’s published LSboost and Factor GDP nowcasts [59]. Table 3 presents forecast-accuracy comparisons based on RMSE and Diebold–Mariano (DM) tests. Moirai Base yields the lowest RMSE, indicating the best point-forecast performance overall. The DM statistics show that Moirai’s errors are significantly smaller than those of a benchmark ARIMA model. Still, it is different from the RBNZ’s LSBoost and Factor models are not statistically significant at conventional levels. Thus, while Moirai demonstrably outperforms ARIMA, it performs on par with LSBoost and Factor an important finding that highlights their comparable predictive accuracy.
@@ 5.1 TSFM Zero-Shot Effectiveness in Macroeconomic Forecasting
Based on RQ1, empirical studies reveal a clear hierarchy among state-of-the-art TSFMs in a zero-shot setting. These TSFMs represent promising steps toward universal forecasting. Empirically, Moirai currently demonstrates the most consistent accuracy on macro and industry series, exhibiting the lowest scaled errors and highest ranks. This is likely attributable to task-specific design and massive pretraining [24]. Chronos also performs strongly, especially on monthly and other high-frequency series matching state-of-the-art zero-shot accuracy with only minor configuration tweaks [23]. By contrast, although TimeGPT offers broad cross-domain versatility, it trails leading alternatives on several specialised economic benchmarks [25]. Both models, however, can significantly narrow this gap when they are fine-tuned or adapted in context with domain-specific economic and industry datasets [66]. Overall, these models generalise better across domains than traditional local models. However, it is important to be aware of their limitations and select model size and type accordingly. Because zero-shot forecasts rely solely on patterns learned during pre-training, they cannot adapt when the data-generating process shifts to a new regime not represented in the training corpus.
@@ 5.2 Effectiveness of TSFMs in Zero-Shot Forecasting during Shocks
5.2.1 Zero-Shot TSFMs vs Domain-Specific Models for Macroeconomic Forecasting
@@ 5.3 Limitations and Future Works
Furthermore, we plan on expanding our future work by moving beyond zero-shot evaluation to explore a fine-tuning approach for retraining models with plausible exogenous shocks. This expansion will also involve extending evaluations to other economies and time periods for accuracy improvement, and applying a quarterly intercept correction to trim forecast errors without compromising cross-domain generality. Lastly, reporting full predictive distributions via density forecasts and coverage metrics for assessing these models deliver calibrated probabilities as well as sharp point forecasts.
@@ 6 Conclusion
This research delivers one of the first systematic zero-shot evaluations of TSFMs for national accounts forecasting. Without any fine-tuning, several TSFM variants, including TimeGPT, Chronos, and Moirai, were tasked with predicting the quarterly year-on-year growth rates of four headline series: National GDP, Primary Industries, Goods-Producing Industries, and Services Industries. Each model operated in a purely univariate setting, eliminating bespoke econometric specifications and long local training histories. Point forecasts and predictive interval calibration are benchmarked against classical/statistical baselines and the RBNZ’s operational models.
The Moirai variants outperformed Persistence and ARIMA across several horizons and match the RBNZ benchmarks. Crucially, TSFMs remained broadly resilient during the COVID-19 structural break, yet the findings reveal small and systematic post-shock biases that suggest a need for calibration with indicator rich models. These results show that carefully engineered TSFMs can internalize rich economic dynamics and deliver actionable forecasts for macroeconomic monitoring and strategic planning. At the same time, clear directions exist for future work targeting, fine-tuning, ensemble blending, and rigorous stress testing across alternative seasonal adjustments and crisis scenarios.

### v2

@@ 3.4 Model Evaluation
In the forecasting literature, scale-dependent error measures remain fundamental for assessing forecast accuracy.
Mean Absolute Error (MAE) and Mean Squared Error/Root Mean Squared Error (MSE/RMSE) in equations 6 and 7 are the two widely reported scale-dependent accuracy metrics. MAE specifically offers a direct interpretation as the average error in the original units of the series [72]. Both compare a forecast $\hat{y}_{t}$ against the observed value $y_{t}$ over $T$ time steps, but they weigh errors differently. MAE is the arithmetic mean of absolute deviations, squares the residuals before averaging, and then takes the square root. But their error weighting differs. MAE averages the absolute errors $|y_{t}-\hat{y}_{t}|$, whereas RMSE averages the squared errors $(y_{t}-\hat{y}_{t})^{2}$ before taking the square root.
@@
MAE makes it easy to interpret the forecast on average.
$\text{MSE}=\frac{1}{T}\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}$ (6)
$\text{RMSE}=\sqrt{\text{MSE}}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}}.$ (7)
MSE/RMSE as defined in equations 6 and 7 squares errors, giving disproportionate weight to large deviations and making the metric sensitive to outliers or significant irregularities. Both share a limitation: they are expressed in the original data units, preventing direct comparison across series with different scales. To address this, Symmetric Mean Absolute Percentage Error (SMAPE) in equation 8 normalises the absolute error by the average magnitude of the actual and forecast values [73],
@@
where $m$ is the seasonal period. This scaling provides a key advantage: MASE is a scale-free error metric. Because the denominator is the MAE of the naive seasonal forecast, the value serves as a direct benchmark. Thus, any MASE $<$ 1 indicates performance superior to the naive baseline, and MASE $>$ 1 denotes inferior accuracy. This scale-free property facilitates widespread adoption in comparative studies of forecasting algorithms.
The Diebold-Mariano (DM) in equations 10, 11, and 12, introduced by [75] established a general, loss-function-agnostic framework for testing whether two competing forecasts have the same expected predictive accuracy.
$e_{i,t}=y_{t}-\hat{y}_{i,t}$ (10)
The forecast error from a model is $i$, a loss function is $\ell(\cdot)$, for example $\ell(e)=e^{2}$ (squared error) or $\ell(e)=|e|$ (absolute error). Define the loss differential:
$d_{t}=\ell\bigl{(}e_{1,t}\bigr{)}-\ell\bigl{(}e_{2,t}\bigr{)}$ (11)
@@
[…]
@@ 3.5 Zero-Shot Forecasts
TSFMs such as TimeGPT, Chronos, and Moirai that employ zero-shot learning represent a significant advance in predictive analytics. Through large-scale pre-training on a wide variety of time series datasets, these models develop a generalised understanding of temporal dynamics, capturing regular cycles and anomalous events. Armed with this broad temporal intuition, they can be deployed directly on novel forecasting tasks without requiring extensive domain-specific fine-tuning. At the core of zero-shot forecasting is the idea that, once a foundation model has internalised patterns spanning many industries and time scales, it can transfer that knowledge seamlessly to new contexts. This approach significantly reduces the time, computational resources, and specialised expertise normally needed for forecasting solutions, as it eliminates the need for extensive re-training and adaptation for each new task.
The practical benefits of zero-shot forecasting are most pronounced in settings where conditions change rapidly or data is scarce. For instance, in sectors such as National GDP, Primary Industries, Goods-Producing Industries, and Services Industries, unexpected events and economic shocks can frequently upend historical relationships and data patterns. In such dynamic environments, a zero-shot model can provide immediate, reasonably accurate projections without waiting for new data to accumulate or models to be re-trained. Furthermore, the same pre-trained model can often be utilised for both short-term operational decisions and longer-term strategic planning, all without needing to rebuild the model for each specific forecasting horizon.
3.6 Experiment Pipeline
@@
We design an experimental back-testing pipeline, shown in Figure 2. That compares shift operation persistence and statistical baseline ARIMA models against zero-shot forecasts from TSFMs: TimeGPT-1, Chronos-T5 (small, base, large), and Moirai (small, base, large). We generate one-quarter-ahead forecasts for each model over an expanding training window from 1999Q3 to 2024Q3. The forecasts are produced without any additional fine-tuning; the models were pre-trained before being evaluated on the same error metrics (MAE, RMSE, SMAPE, and MASE). To capture the heterogeneous dynamics of New Zealand’s economy, we apply this experiment to four broad production sectors: National GDP, Primary Industries, Goods-Producing Industries, and Services Industries. The selection of these models for the experiments also reveals which classical and TSFMs are strongest with scarce data.
@@ 4.2 Forecast Benchmarking Against State-of-the-art
Figure 5: Comparisons of the Moirai Base Model with LSboost and Factor Model using Reserve Bank Nowcasts.
We extracted forecasts from the RBNZ’s LSboost (Gradient boosting) and Factor model, observing that they are quarterly percentage changes. Consequently, we converted the national GDP data from a quarterly annual percentage change to a quarterly percentage change to match the forecast format. For comparison in Figure 5, we also used Moirai Base to forecast quarterly percentage change values. The table includes Diebold Mariano two-sided tests computed against an ARIMA model for the period 2009Q1 - 2019Q1, evaluating the RBNZ’s published LSboost and Factor GDP nowcasts [66]. Table 3 presents forecast accuracy comparisons based on RMSE and Diebold–Mariano (DM) tests. Moirai Base yields the lowest RMSE, indicating the best point-forecast performance overall. The DM statistics show that Moirai’s errors are significantly smaller than those of a benchmark ARIMA model. Still, the differences from the RBNZ’s LSBoost and Factor models are not statistically significant at conventional levels. Thus, while Moirai demonstrably outperforms ARIMA, it performs on par with LSBoost and Factor, an important finding that highlights their comparable predictive accuracy.
@@ 5.1 TSFM Zero-Shot Effectiveness in Macroeconomic Forecasting
Based on RQ1, empirical studies reveal a clear hierarchy among state-of-the-art TSFMs in a zero-shot setting. These TSFMs represent promising steps toward universal forecasting. Empirically, Moirai currently demonstrates the most consistent accuracy on macro and industry series, exhibiting the lowest scaled errors and highest ranks. This is likely attributable to task-specific design and massive pre-training [25]. Chronos also performs strongly, especially on monthly and other high-frequency series matching state-of-the-art zero-shot accuracy with only minor configuration tweaks [26]. By contrast, although TimeGPT offers broad cross-domain versatility, it trails leading alternatives on several specialised economic benchmarks [24]. Both models, however, can significantly narrow this gap when they are fine-tuned or adapted in context with domain-specific economic and industry datasets [76]. Overall, these models generalise better across domains than traditional local models. However, it is important to be aware of their limitations and select model size and type accordingly. Because zero-shot forecasts rely solely on patterns learned during pre-training, they cannot adapt when the data-generating process shifts to a new regime not represented in the training corpus.
@@ 5.2 Effectiveness of TSFMs in Zero-Shot Forecasting during Shocks
5.3 Zero-Shot TSFMs vs Domain-Specific Models for Macroeconomic Forecasting
@@ 5.4 Limitations and Future Works
Furthermore, we plan on expanding our future work by moving beyond zero-shot evaluation to explore a fine-tuning approach for re-training models with plausible exogenous shocks. This expansion will also involve extending evaluations to other economies and time periods for accuracy improvement, and applying a quarterly intercept correction to trim forecast errors without compromising cross-domain generality. Lastly, reporting full predictive distributions via density forecasts and coverage metrics for assessing these models deliver calibrated probabilities as well as sharp point forecasts.
@@ 6 Conclusion
This research delivers one of the first systematic zero-shot evaluations of TSFMs for macroeconomic indicators forecasting. Without any fine-tuning, several TSFM variants, including TimeGPT, Chronos, and Moirai, were tasked with predicting the quarterly year-on-year growth rates of four headline series: National GDP, Primary Industries, Goods-Producing Industries, and Services Industries. Each model operated in a purely univariate setting, eliminating bespoke econometric specifications and long local training histories. Point forecasts and predictive interval calibration were benchmarked against classical/statistical baselines and operationalised models from large institutions.
The Moirai variants outperformed Persistence and ARIMA across several horizons and matched industry benchmarks. Crucially, TSFMs remained broadly resilient during the COVID-19 structural break, yet the findings revealed small and systematic post-shock biases that suggest a need for multivariate extensions that incorporate other indicators. These results show that carefully engineered TSFMs can internalise rich economic dynamics and deliver actionable forecasts for macroeconomic monitoring and strategic planning. At the same time, clear directions exist for future work targeting, multivariate integrations, fine-tuning, ensemble blending, and rigorous stress testing across additional crisis scenarios.